# Adaptive Massively Parallel Coloring in Sparse Graphs

### v1

@@ 6.1 An $O(\alpha^{2+\varepsilon})$-coloring in $O(1/\varepsilon)$ Rounds
Compute a $\beta$-partition using Theorem 1.2 with $\beta=\alpha^{1+\varepsilon}$ for a constant $\varepsilon>0$. This will require $O(1/\varepsilon)$ AMPC rounds and the $\beta$-partition will have $O(\varepsilon^{-1}\log_{\alpha}n)$ layers.
@@
Note that if $\alpha^{\varepsilon}>\log n$, we don’t need to simulate all $\log^{*}n$ rounds of the Arb-Linial algorithm. If we just simulate the first round of Arb-Linial, we get an $O(\beta^{2}\log n)$-coloring. Since $\beta=\alpha^{1+\varepsilon}$ and $\log n<\alpha^{\varepsilon}$, $O(\beta^{2}\log n)=O(\alpha^{2+3\varepsilon})$, which we can say is $O(\alpha^{2+\varepsilon})$, by replacing $\varepsilon$ by $\varepsilon/3$. We can simulate the first round of Arb-Linial in $1$ AMPC round by gathering the at most $\beta$ out-neighbors of each node. We can do this since we assumed $\alpha\leq n^{\delta/(1+\varepsilon)}$, which implies $\beta\leq n^{\delta}$.
@@ 6.4 Coloring when $\alpha$ is Large
We will prove this theorem by providing a deterministic, non-component-stable, MPC algorithm that produces the coloring in $O(\log_{x}n)$ rounds using $O(n+m)\operatorname{\text{{\rm poly}}}\log n$ global memory. Since the AMPC model can simulate any MPC algorithm without any loss in round complexity, the theorem follows. Consider the following randomized trial with pairwise independent random coins: Every node picks a random color from $1,\dots,2x\Delta$. Let $Y_{e}$ be an indicator random variable for the event that edge $e$ is monochromatic, and let $Y=\sum_{e}Y_{e}$ be the number of monochromatic edges produced by the randomized trial. Since we use pairwise independent random coins, the probability that $Y_{e}=1$ is $1/(2x\Delta)$. By the handshaking lemma and linearity of expectation, $\operatorname*{\mathbf{E}}[Y]\leq\sum_{v}\sum_{e\ni v}\operatorname*{\mathbf{E}}[Y_{e}]=\sum_{v}\deg(v)/(2x\Delta)\leq|V|/(2x)$. We derandomize this process using the method of conditional expectations, which was first introduced in the context of the MPC model in [CPS20]. Since we use pairwise independence, we can reduce the number of random coins required to $O(\log n)$-bits by using a family of pairwise independent hash functions. Then we fix these random bits in batches of size $\lfloor(\delta/3)\log_{2}n\rfloor$, and we denote the $i^{th}$ batch by $B_{i}$. The invariant we will maintain when fixing the random bits in each batch is that $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]\leq\operatorname*{\mathbf{E}}[Y]$. When fixing batch $B_{i}$, let us assume inductively that the invariant is maintained for all previous batches, i.e. we have fixed the random bits such that $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i-1}=b_{i-1}]\leq\operatorname*{\mathbf{E}}[Y]$. By the law of total expectations, there exists an assignment $B_{i}=b_{i}$ such that the invariant is maintained. There are at most $n^{\delta/3}$ assignments to $B_{i}$, and we will calculate the value of the conditional expectation $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]$ for each assignment $b_{i}$ in parallel and pick the minimum. Let us see first how to calculate the conditional expectation for a single assignment $B_{i}=b_{i}$. The key idea is that we decompose $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]$ into $\sum_{e}\operatorname*{\mathbf{E}}[Y_{e}|B_{1}=b_{1},\dots,B_{i}=b_{i}]$, so the task is to compute a sum of values where each value can be computed locally by the machine that stores edge $e$. We create a “broadcast tree” on the set of machines, which is a balanced $n^{\delta/2}$-ary tree of depth $O(1/\delta)$. Each machine $M$ locally computes $\operatorname*{\mathbf{E}}[Y_{e}|B_{1}=b_{1},\dots,B_{i}=b_{i}]$ for all edges $e$ stored at $M$ and adds them up. The leaves in the broadcast tree simply send this sum to their parent. Once a machine in the broadcast tree receives the values from all its children, it sends the sum of all these values and its locally computed sum to its parent. In each round, a machine sends at most one value and receives at most $n^{\delta/2}$ values. In $O(1/\delta)$ rounds, the root of the broadcast tree will calculate the value of the conditional expectation $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]$. Since there are at most $n^{\delta/3}$ assignments to $B_{i}$, it is easy to see that the machines can calculate the conditional expectation value for all assignments $b_{i}$ in parallel over the same broadcast tree. In each round, a machine sends at most $n^{\delta/3}$ values and receives at most $n^{5\delta/6}$ values ($n^{\delta/3}$ values from each of its $n^{\delta/2}$ children in the broadcast tree). Therefore the root of the tree can pick the assignment $b_{i}$ which minimizes the conditional expectation, and this maintains the invariant for our assignment to $B_{i}$. Since there are $O(1/\delta)$ batches, we are able to fix all the random bits in $O(1/\delta^{2})$ rounds. Now we have a deterministic algorithm that colors nodes with colors $1,\dots,2x\Delta$, such that there are at most $n/(2x)$ monochromatic edges. We fix the color of all nodes that do not have an incident monochromatic edge and all other nodes are uncolored. Let $U$ be the set of uncolored nodes. We have that $|U|\leq n/x$. If we repeat the same randomized process for the uncolored nodes, i.e., each node in $U$ picks a color in $1,\dots,2x\Delta$ using pairwise independent random bits. We still have that the expected number of monochromatic edges is $\operatorname*{\mathbf{E}}[Y]\leq|U|/(2x)\leq n/(2x^{2})$. We again derandomize this process, and we get a coloring of nodes with colors $1,\dots,2x\Delta$, such that there are at most $n/(2x^{2})$ monochromatic edges. The new set of uncolored nodes has size at most $n/x^{2}$. Therefore, if we repeat the process $i$ times we get a partial coloring of nodes with colors $1,\dots,2x\Delta$, such that the number of uncolored nodes is at most $n/x^{i}$. We continue until no uncolored vertices remain. The number of repetitions is at most $O(\log_{x}n)$, and we use at most $2x\Delta$ colors. ∎

### v2

@@ 6.1 An $O(\alpha^{2+\varepsilon})$-coloring in $O(1/\varepsilon)$ Rounds
Compute a $\beta$-partition using Theorem 1.2 with $\beta=O(\alpha^{1+\varepsilon})$ for a constant $\varepsilon>0$. This will require $O(1/\varepsilon)$ AMPC rounds and the $\beta$-partition will have $O(\varepsilon^{-1}\log_{\alpha}n)$ layers.
@@
Note that if $\alpha^{\varepsilon}>\log n$, we don’t need to simulate all $\log^{*}n$ rounds of the Arb-Linial algorithm. If we just simulate the first round of Arb-Linial, we get an $O(\beta^{2}\log n)$-coloring. Since $\beta=O(\alpha^{1+\varepsilon})$ and $\log n<\alpha^{\varepsilon}$, $O(\beta^{2}\log n)=O(\alpha^{2+3\varepsilon})$, which we can say is $O(\alpha^{2+\varepsilon})$, by replacing $\varepsilon$ by $\varepsilon/3$. We can simulate the first round of Arb-Linial in $1$ AMPC round by gathering the at most $\beta$ out-neighbors of each node. We can do this since we assumed $\alpha\leq n^{\delta/(1+\varepsilon)}$, which implies $\beta\leq n^{\delta}$.
@@ 6.4 Coloring when $\alpha$ is Large
We will prove this theorem by providing a deterministic, non-component-stable, MPC algorithm that produces the coloring in $O(\log_{x}n)$ rounds using $O(n+m)\operatorname{\text{{\rm poly}}}\log n$ global memory. Since the AMPC model can simulate any MPC algorithm without any loss in round complexity, the theorem follows.
Consider the following randomized trial with pairwise independent random coins: Every node picks a random color from $1,\dots,2x\Delta$. Let $Y_{e}$ be an indicator random variable for the event that edge $e$ is monochromatic, and let $Y=\sum_{e}Y_{e}$ be the number of monochromatic edges produced by the randomized trial. Since we use pairwise independent random coins, the probability that $Y_{e}=1$ is $1/(2x\Delta)$. By the handshaking lemma and linearity of expectation, $\operatorname*{\mathbf{E}}[Y]\leq\sum_{v}\sum_{e\ni v}\operatorname*{\mathbf{E}}[Y_{e}]=\sum_{v}\deg(v)/(2x\Delta)\leq|V|/(2x)$.
We derandomize this process using the method of conditional expectations, which was first introduced in the context of the MPC model in [CPS20]. Since we use pairwise independence, we can reduce the number of random coins required to $O(\log n)$-bits by using a family of pairwise independent hash functions. Then we fix these random bits in batches of size $\lfloor(\delta/3)\log_{2}n\rfloor$, and we denote the $i^{th}$ batch by $B_{i}$. The invariant we will maintain when fixing the random bits in each batch is that $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]\leq\operatorname*{\mathbf{E}}[Y]$.
When fixing batch $B_{i}$, let us assume inductively that the invariant is maintained for all previous batches, i.e. we have fixed the random bits such that $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i-1}=b_{i-1}]\leq\operatorname*{\mathbf{E}}[Y]$. By the law of total expectations, there exists an assignment $B_{i}=b_{i}$ such that the invariant is maintained. There are at most $n^{\delta/3}$ assignments to $B_{i}$, and we will calculate the value of the conditional expectation $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]$ for each assignment $b_{i}$ in parallel and pick the minimum.
Let us see first how to calculate the conditional expectation for a single assignment $B_{i}=b_{i}$. The key idea is that we decompose $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]$ into $\sum_{e}\operatorname*{\mathbf{E}}[Y_{e}|B_{1}=b_{1},\dots,B_{i}=b_{i}]$, so the task is to compute a sum of values where each value can be computed locally by the machine that stores edge $e$. We create a “broadcast tree” on the set of machines, which is a balanced $n^{\delta/2}$-ary tree of depth $O(1/\delta)$. Each machine $M$ locally computes $\operatorname*{\mathbf{E}}[Y_{e}|B_{1}=b_{1},\dots,B_{i}=b_{i}]$ for all edges $e$ stored at $M$ and adds them up. The leaves in the broadcast tree simply send this sum to their parent. Once a machine in the broadcast tree receives the values from all its children, it sends the sum of all these values and its locally computed sum to its parent. In each round, a machine sends at most one value and receives at most $n^{\delta/2}$ values. In $O(1/\delta)$ rounds, the root of the broadcast tree will calculate the value of the conditional expectation $\operatorname*{\mathbf{E}}[Y|B_{1}=b_{1},\dots,B_{i}=b_{i}]$.
Since there are at most $n^{\delta/3}$ assignments to $B_{i}$, it is easy to see that the machines can calculate the conditional expectation value for all assignments $b_{i}$ in parallel over the same broadcast tree. In each round, a machine sends at most $n^{\delta/3}$ values and receives at most $n^{5\delta/6}$ values ($n^{\delta/3}$ values from each of its $n^{\delta/2}$ children in the broadcast tree). Therefore the root of the tree can pick the assignment $b_{i}$ which minimizes the conditional expectation, and this maintains the invariant for our assignment to $B_{i}$. Since there are $O(1/\delta)$ batches, we are able to fix all the random bits in $O(1/\delta^{2})$ rounds.
Now we have a deterministic algorithm that colors nodes with colors $1,\dots,2x\Delta$, such that there are at most $n/(2x)$ monochromatic edges. We fix the color of all nodes that do not have an incident monochromatic edge and all other nodes are uncolored. Let $U$ be the set of uncolored nodes. We have that $|U|\leq n/x$. If we repeat the same randomized process for the uncolored nodes, i.e., each node in $U$ picks a color in $1,\dots,2x\Delta$ using pairwise independent random bits. We still have that the expected number of monochromatic edges is $\operatorname*{\mathbf{E}}[Y]\leq|U|/(2x)\leq n/(2x^{2})$. We again derandomize this process, and we get a coloring of nodes with colors $1,\dots,2x\Delta$, such that there are at most $n/(2x^{2})$ monochromatic edges. The new set of uncolored nodes has size at most $n/x^{2}$. Therefore, if we repeat the process $i$ times we get a partial coloring of nodes with colors $1,\dots,2x\Delta$, such that the number of uncolored nodes is at most $n/x^{i}$. We continue until no uncolored vertices remain. The number of repetitions is at most $O(\log_{x}n)$, and we use at most $2x\Delta$ colors. ∎